# GaussianRoom: Improving 3D Gaussian Splatting with SDF Guidance and Monocular Cues for Indoor Scene Reconstruction

### v1

@@ 5.1 Experimental Setup
5.2 Reconstruction Evaluation
Tab. 1 shows our method significantly outperforms both Gaussian-based methods and NeRF-based methods in geometry metrics on the ScanNet and ScanNet++ datasets. As illustrated in Fig. 3, Gaussian-based scene reconstruction methods, such as SuGaR [18] and DN-Splatter [48], are impacted by the geometric disorder of Gaussians, leading to uneven Poisson reconstruction results. In contrast, our methods can achieve smooth and continuous results due to the continuity of neural SDF. Compared with NeRF-based methods, GaussianRoom achieves more complete and detailed reconstruction while greatly shortening training time, due to integrated 3DGS encouraging more efficient point sampling near the surface and edge prior offers more attention to detailed regions. For example, MonoSDF requires about 18-hour optimization, whereas GaussianRoom achieves better results within 4-hour training. Please refer to the supplementary for more qualitative results on ScanNet++ [59] dataset compared with existing methods.
@@ 5.2 Reconstruction Evaluation
Table 1: Quantitative reconstruction comparison on ScanNet / ScanNet++ datasets. We report the average results over 8 and 2 scenes respectively. The best results are marked in bold.
@@
Table 2: Quantitative rendering comparison with existing methods on ScanNet and ScanNet++ datasets. We report the average results over 8 and 2 scenes respectively.
@@
Figure 4: Qualitative rendering comparisons. As shown from the above-highlighted patches, the rendering results of GaussianRoom significantly outperform other GS-based methods, including texture-less regions and details.
5.3 Rendering Evaluation
@@
According to Tab. 2, our method achieves superior rendering metrics on both the ScanNet and ScanNet++ datasets compared to 3DGS [24], SuGaR [18], GaussianPro [9], and DN-Splatter [48]. The substantial improvements in the PSNR metric indicate that our rendered images exhibit significantly reduced distortion, credited to our designed SDF-guided Primitive Distribution Strategy, which effectively prunes floating Gaussian primitives in non-surface space and deploys Gaussians near the surface that lacks initial Gaussians, as demonstrated in Fig. 2 (c). As shown in Fig. 4, our method achieves sharp details and superior renderings in both rich-texture and low-texture regions, benefiting from our rational utilization of the geometric information within the neural SDF to align Gaussian primitives with the surface of the scene.
5.4 Ablation Study
In this section, we conduct experiments by individually removing each improvement module from our full model to verify the effectiveness of each design. Quantitative results and qualitative visualizations can be found in Tab. 3 and Fig. 5.
3D Reconstruction Novel View Synthesis
Method Accuracy $\downarrow$ Completion $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ F-score $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
w/o SGD 0.048 0.044 0.790 0.730 0.758 0.756 23.545 0.394
w/o SDP 0.048 0.044 0.785 0.724 0.753 0.738 22.589 0.395
w/o Gaussian-guided sampling 0.049 0.045 0.777 0.718 0.746 0.754 23.463 0.395
w/o normal prior 0.127 0.120 0.412 0.367 0.388 0.755 23.625 0.395
w/o edge prior 0.055 0.046 0.780 0.722 0.750 0.754 23.350 0.397
GaussianRoom (Full) 0.047 0.043 0.800 0.739 0.768 0.758 23.601 0.391
@@ 5.4 Ablation Study
Table 3: Ablation study in the proposed modules. The ablation experiments were conducted on 8 scenes from the ScanNet dataset, with the best metrics highlighted in bold.
Figure 5: Ablation studies. The above highlighted patches visually show the rendering degradation caused by each missing module.
The experimental results demonstrate that our SDF-guided Densification and Pruning (SDP) and SDF-guided Global Densification (SGD) modules effectively impose geometric constraints on disordered Gaussians, thereby enhancing rendering quality and providing more accurate depth information. With the depth information, improvements can be made to the sampling precision and convergence speed of the neural SDF through the Gaussian-guided Point Sampling (GPS) module, creating a mutually beneficial feedback loop. The ablation studies of the normal prior and edge prior demonstrate that these two components respectively enhance 3D reconstruction quality in low-texture regions and detailed regions. Given the large low-texture areas typical in indoor reconstruction tasks, the supervision provided by the normal prior proves highly effective.
@@
However, similar to DN-Splatter [48], utilizing the normal prior as a regularization term to constrain the orientation of Gaussians results in a decrease in the PSNR metric. Despite this, our full model excels in both SSIM and LPIPS metrics, achieving superior visualization quality in detailed regions, as depicted in Fig. 5.
6 Conclusion
@@
Appendix A Appendix / supplemental material
The supplementary material section is organized as follows: 1) The first section elaborates on the implementation details of GaussianRoom, covering hyper-parameters and geometry evaluation details. 2) We then show additional experimental results. 3) Lastly, we delve into limitations and future directions.
A.1 Implementation Details
During the co-optimization stage of 3DGS and neural SDF, we implement the SDF-guided Densification and Pruning strategy every 100 iterations and the SDF-guided Global Densification strategy every 2000 iterations, rather than employing the original Gaussian densification and pruning strategy, with other hyper-parameters remaining mostly consistent with 3DGS. As for the neural SDF, We sample 1024 rays per batch and 64+64 points on each ray. After training, we extract a mesh from the SDF by the Marching Cube algorithm [34] with the volume size of $512^{3}$. For geometry comparison, we utilize Poisson reconstruction to get the mesh for COLMAP [42] and some Gaussian-based methods [24, 9, 48] and set octree depth = 11.
Below, we introduce the hyper-parameters used in our experiments:
  • For the edge-aware neural SDF optimization strategy, the hyperparameter $\delta$ indicating the importance of edges is set to 2.
  • For the resolution of partitioned scene, we set $\mathrm{N}=128$.
  • For the loss weights discussed in Sec. 4.3, we set $\lambda_{1}$ = 0.8, $\lambda_{2}$ = 0.01 and $\lambda_{eik}$ = 0.1.
  • For taking K nearest neighbor Gaussians to initialize the new Gaussians, we set $K$ = 10
  • We execute the SDF-guided Densification and Pruning strategy every $F_{SDP}$ iterations, and we set $F_{SDP}$ = 100
  • We execute the SDF-guided Global Densification strategy every $F_{SGD}$ iterations, and we set $F_{SGD}$ = 1999
  • The other hyperparameters used in the experiment are as follows: $\tau_{s}$ = 0.01, $\tau_{n}$ = 10, $\tau_{p}$ = 0.998, $\tau_{d}$ = 0.999, $\tau_{g}$ = 0.0002, $\gamma$ = 5.
[…]
where $f_{g}(\cdot)$ represents the geometry network of neural SDF.
Evaluation Details For 3D geometry evaluation, the output mesh by some methods [52, 51, 62] would reconstruct some areas out of the scope of the ground truth. So the direct evaluation between them is not fair. To address the above issues, similar to [51], we first remove the faces that cannot be observed by any input camera pose. Then we filter faces at empty areas of the ground truth using 2D mask information rendered by the ground truth for each input view. Tab. 4 define the 3D geometry metrics for evaluation of the main paper. Among these metrics, F-score is considered as the most representative metric to evaluate geometry quality, which contains the information of both accuracy and completeness.
3D Metics
Acc. $\mbox{mean}_{p\in P}(\min_{p^{*}\in P^{*}}||p-p^{*}||)$
Comp. $\mbox{mean}_{p^{*}\in P^{*}}(\min_{p\in P}||p-p^{*}||)$
[…]
F-score $\frac{2\times\text{Perc}\times\text{Recall}}{\text{Perc}+\text{Recall}}$
Table 4: Definitions of 3D geometry metrics: $p$ and $p^{*}$ are the predicted and ground truth depths.
A.2 More Visualization Results
In this section, we give more visualization results of our method compared with other methods.
Rendering Comparisons We compare visualize rendering results with 3DGS [24], SuGaR [18] and DN-Splatter [48]. As shown in Fig. 6 and Fig. 7, GaussianRoom significantly outperforms other GS-based methods in the highlighted patches, including texture-less regions and details.
Figure 6: Qualitative rendering comparisons on ScanNet dataset. As shown from the above highlighted patches, the rendering results of GaussianRoom significantly outperform other GS-based methods, including texture-less regions and details.
Figure 7: Qualitative rendering comparisons on ScanNet and ScanNet++ datasets. As shown from the above highlighted patches, the rendering results of GaussianRoom significantly outperform other GS-based methods, including texture-less regions and details.
Reconstruction Comparisons For ScanNet++ dataset, we compare with NeRF [25], NeuS [52], MonoSDF [62], SuGaR [18] and DN-Splatter [48]. As shown in Fig. 8, our method can reconstruct finer details of indoor scenes.
Figure 8: Qualitative reconstruction comparisons on ScanNet++. For each indoor scene, the first row is the top view of the whole room, and the second row is the details of the masked region.
A.3 Limitations and Future Works
Although the training time of our method is greatly improved compared to existing NeRF-based reconstruction methods, the efficiency of our neural SDF still lags behind the 3DGS, leading to extended training time compared to 3DGS only. Hence, improving the efficiency of the MLP-based neural SDF is a promising direction for our future work.

### v2

@@ V-A Experimental Setup
V-B Reconstruction Evaluation
Tab. I shows GaussianRoom outperforms both Gaussian-based and NeRF-based methods in geometry metrics on the ScanNet and ScanNet++ datasets. As illustrated in Fig. 4, Gaussian-based scene reconstruction methods, such as SuGaR [14] and DN-Splatter [44], are impacted by the geometric disorder of Gaussians, leading to uneven Poisson reconstructions. In contrast, our methods can achieve smooth and continuous results due to the continuity of neural SDF. Compared with NeRF-based methods, GaussianRoom achieves more complete and detailed reconstruction while greatly shortening training time, due to integrated 3DGS encouraging more efficient point sampling near the surface and edge prior offers more attention to detailed regions. For example, MonoSDF requires about 18-hour optimization, whereas GaussianRoom achieves better results within 5-hour training.
TABLE I: Quantitative reconstruction comparison on ScanNet / ScanNet++. Average results are reported over 8 and 2 scenes, respectively, with the best results in bold.
@@ V-B Reconstruction Evaluation
TABLE II: Quantitative rendering comparison with existing methods on ScanNet and ScanNet++ datasets. We report the average results over 8 and 2 scenes respectively.
@@
Figure 5: Qualitative rendering comparisons. As shown from the above-highlighted patches, the rendering results of GaussianRoom outperform other GS-based methods, including texture-less regions and details.
V-C Rendering Evaluation
@@
According to Tab. II, our method demonstrates higher rendering metrics on both the ScanNet and ScanNet++ datasets compared to 3DGS [22], SuGaR [14], GaussianPro [6], and DN-Splatter [44]. The substantial improvements in the PSNR metric indicate that our rendered images exhibit reduced distortion, credited to our designed SDF-guided Primitive Distribution Strategy, which effectively prunes floating Gaussian primitives in non-surface space and deploys Gaussians near the surface that lacks initial Gaussians, as demonstrated in Fig. 3. As shown in Fig. 5, our method achieves sharp details and improved renderings in both rich-texture and low-texture areas by leveraging geometric information in the neural SDF to align Gaussian primitives with the scene surface.
V-D Ablation Study
In this section, we conduct ablation experiments to verify the superiority of mutual learning of 3DGS and SDF fields and the effectiveness of monocular cues. As shown in Tab. III, training neural SDF is time-consuming and its rendering quality and speed are inferior compared to 3DGS. In contrast, 3DGS struggles to learn geometry accurately, resulting in poor reconstruction quality. Compared to them, GaussianRoom not only avoids their weaknesses but also brings further improvements in both reconstruction and rendering quality. GaussianRoom requires only half the training time of SDF while achieving improved reconstruction results. Additionally, its rendering quality surpasses that of 3DGS with monocular cues, while maintaining a rendering speed of over 170 fps.
@@ V-D Ablation Study
TABLE III: Ablation study of mutual optimization. The ablation experiments were conducted on 8 scenes from the ScanNet dataset, with the best metrics highlighted in bold.
3D Reconstruction Novel View Systhesis Rendering speed
Method F-score $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ fps $\uparrow$ Training time $\downarrow$
GaussianRoom(full) 0.768 0.758 23.601 0.391 170+ $\sim$4.5h
SDF + MonoCues 0.719 0.714 22.634 0.458 $<$1 $\sim$9h
3DGS + MonoCues 0.163 0.726 22.092 0.396 170+ $\sim$1h
w/o normal prior 0.388 0.755 23.625 0.395 170+ $\sim$4.5h
w/o edge prior 0.750 0.754 23.35 0.397 170+ $\sim$4.5h
@@
Ablation studies in Tab. III also show that the normal prior and edge prior improve 3D reconstruction quality, with each primarily functioning in low-texture regions and detailed regions, respectively. Given the large low-texture areas typical in indoor reconstruction tasks, the supervision provided by the normal prior proves to be highly effective. However, similar to DN-Splatter [44], utilizing the normal prior as a regularization term to constrain the orientation of Gaussians results in a decrease in the PSNR metric. Despite this, our full model excels in both SSIM and LPIPS metrics.
VI CONCLUSIONS
@@
ACKNOWLEDGMENT
This research was funded by Shenzhen Major Science and Technology Project (KJZD20230923115503007)